In isosceles triangle CDE, angle C is thirty-nine degrees more than angle D. angle C is one of the congruent angles. What is the measure of each angle in triangle CDE?
Answer by stanbon(75887)

In isosceles triangle CDE, angle C is thirty-nine degrees more than angle D. angle C is one of the congruent angles. What is the measure of each angle in triangle CDE?
-------------------
Let the vertex angle be "x" degrees.
Each base angle is "x+39" degrees.
-----------
EQUATION:
x + 2(x+39) = 180
3x + 78 = 180
3x = 102
x = 34 degrees (vertex angle measure)
x+39 = 73 (each base angle meansure)
===============================
Cheers,
Stan H.